OpenStudy (anonymous):

please help thank you so much loves. Identify the x-intercept of g(x). f(x) = (x – 4)^2 g(x) = f(x – 5)

OpenStudy (anonymous):

g(x)=f(x-5)=((x-5)-4)^2 = (x-9)^2 g(x) = (x-9)^2, the x-intercept will be the root of g(x)

OpenStudy (anonymous):

The root of a function g(x) is a value of x such that g(x) = 0

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

no I am confused

OpenStudy (anonymous):

But did u understand how i found g(x) = (x-9)^2 ?

OpenStudy (anonymous):

oh ok yes.

OpenStudy (anonymous):

Okay, now you wnat to find the x-intercept, any function will intercept the x-axis when the functions equals to zero, that is g(x) = 0, so... g(x) = (x-9)^2 = 0, now just solve for x and that is your x-intercept

OpenStudy (anonymous):

would it be 9

OpenStudy (anonymous):

That's right!

OpenStudy (anonymous):

thank you!